An account pays 5% interest, compounded continuously. How long will it take for the investment to double?
Answer by stanbon(75887)

An account pays 5% interest, compounded continuously. How long will it take for the investment to double?
----------------
A(t) = P*e^(rt)
-----
Your Problem:
2P = P*e^(0.05t)
-----
e^(0.05t) = 2
----
0.05t = ln(2)
---
t = (ln(2))/0.05
----
t = 13.86 years
----
It doubles in the 14th year.
------------
Cheers,
Stan H.
--------------------